# MEGConformer: Conformer-Based MEG Decoder for Robust Speech and Phoneme Classification

### v1

@@ 6 Discussion and conclusion
As shown in Table 3, dynamic grouping (vs. fixed groups) yields the largest improvement (+13.3% relative; $W=1.0$, $p=0.004$). Inverse–square‐root class weighting (our default) outperforms non-weighted loss by +7.6% (not significant with $W=10.0$, $p=0.084$). The custom Conformer provides a small lift over Conformer Small (+0.5%; $W=16.0$, $p=0.844$), consistent with similar capacity but with a better fit to the data characteristics. As noted in the main text, holdout ensembling adds +15.4% but is specific to the competition evaluation.
Appendix Appendix B RMS energy analysis across splits in Phoneme Task
To quantify the scale mismatch between partitions, we computed the per-window root-mean-square (RMS) energy across channels and time as in Equation 2, where $x$ is the MEG window (C=306C{=}306 channels, T=125T{=}125 samples). For validation and test, windows were read with the standard competition preprocessing and standardization enabled. The holdout was read as stored (already standardized on disk). We then formed outline-only histograms using the same bin edges and density normalization across splits.
@@
Appendix Appendix C Phonetic feature decoding challenges
To explore linguistically grounded representations for MEG decoding, we trained separate binary classifiers targeting core phonetic features related to the manner of articulation and voicing without sample averaging. Each model was adapted from the Phoneme Classification configuration described in Section 3.3, using the same Conformer backbone and training hyperparameters, but modified for binary output. Notably, we replaced the 39-class head with a single-logit output, used a binary cross-entropy loss with label smoothing and automatic positive-class weighting following Equation 1 to address class imbalance (based on the Speech Detection model in Section 3.2).
This approach follows the ideas proposed by the PNPL competition authors in their blog on language-inspired phoneme classification [kwon2025], where phonetic features are suggested to handle dataset imbalance and improve infrequent phoneme decoding.
Based on [mesgarani2014], we defined six phonetic features based on manner of articulation: Plosive, Fricative, Affricate, Nasal, Liquid, Glide, and a Voicing feature that distinguishes voiced from voiceless phonemes, including vowels and consonants. Table 4 reports preliminary decoding results using F1-macro, positive-class F1, balanced accuracy, and AUROC (mean $\pm$ std across ten seeds).
Following Appendix A, statistical significance is evaluated with a one-sample Wilcoxon signed-rank test against chance ($0.5$ for F1-macro).
@@
Among these features, Voicing showed the strongest decoding signal, approaching 58% F1-macro (significant with $W=55.0$, $p=0.001$). Despite its similarity in label balance to the Speech Task, scores are not on par, implying that decoding some fine-grained speech features may not be straightforward. Manner-based features such as Plosive and Fricative were moderately decodable (both significant with $W=55.0$, $p=0.001$), while Affricate, a composite and low-frequency category (0.97%), remained at chance (not significant with $W=0.0$, $p=1.0$). Attempts to address the affricate issue through ensembling of Plosive and Fricative models, transfer learning from these features, and larger effective batch sizes (via gradient accumulation) did not yield convergence improvements. This supports the view that decomposing phonemes into binary linguistic features alone is insufficient to overcome the data scarcity, as some linguistic features may also be rare enough not to be easily decodable, like affricates here.
@@
Overall, these experiments suggest that feature-based representations are promising for improving interpretability and possibly scaling, but further research is needed to address low-frequency classes and explore multitask formulations that jointly learn shared articulatory subspaces.
Appendix Appendix D Data-size ablation
To examine the effect of dataset scale on decoding performance, we conducted data-size ablations for both tasks using progressively larger subsets of the LibriBrain training data. Figures 4 and 5 show the resulting F1-macro scores as a function of the total hours of MEG data used for training, with shaded bands indicating one standard deviation across random seeds.
Figure 4: Data size ablation for Speech Task.
Figure 5: Data size ablation for Phoneme Task.
For Speech Detection, performance rapidly improved with additional data and then began to saturate, suggesting that the task may already approach its ceiling with a single subject’s data. In contrast, Phoneme Classification, even though with some diminishing returns signs, shows an upward trend with no apparent plateau, implying that larger-scale within-subject recordings could still yield substantial gains. These results align with prior observations [ozdogan2025] that decoding performance continues to scale with the amount of high-quality MEG data available for training.

### v2

@@ 7.1.2 Phoneme task ablation
As shown in Table 3, dynamic grouping (vs. fixed groups) gives the largest improvement (+13.3% relative; $W=1.0$, $p=0.004$). Inverse-square-root class weighting (our default) outperforms non-weighted loss by +7.6% (not significant with $W=10.0$, $p=0.084$). The custom Conformer provides a small lift over Conformer Small (+0.5%; $W=16.0$, $p=0.844$), consistent with similar capacity but with a better fit to the data characteristics.
7.2 Data-size ablation
To assess how the performance scales with within-subject MEG data, we conducted data-size ablations for both tasks using progressively larger subsets of the LibriBrain training set. Figures 4 and 5 show the resulting F1-macro scores as a function of the total hours of MEG data used for training (for phonemes, without sample averaging), with shaded bands indicating one standard deviation across random seeds.
For Speech Detection, performance rapidly improved with additional data and then began to saturate, suggesting that the task may already approach its ceiling with a single subject's data. In contrast, Phoneme Classification, even though with some diminishing returns signs [kaplan2020], shows an upward trend with no apparent plateau, implying that larger-scale within-subject recordings could still give substantial gains. These results align with prior observations [ozdogan2025] that decoding performance continues to scale with the amount of high-quality MEG data available for training.
Figure 4: Data size ablation for Speech Detection Task.
Figure 5: Data size ablation for Phoneme Classification Task, trained without sample averaging
7.3 Frequency band contribution to the tasks
[…]
Overall, both tasks emphasize low-frequency components, with negligible standalone contribution from Gamma and HGA in this dataset. This pattern is consistent with MEG evidence that low-frequency (Delta/Theta) activity supports speech decoding [cogan2011], whereas higher-frequency information appears more accessible in invasive recordings [proix2022], and also aligns with earlier frequency-based MEG decoding analyses [dezuazo2025].
Figure 6: Decoding mean scores by frequency band for Speech and Phoneme Tasks (left/right y-axes), with black bars showing their standard deviation. The dashed lines represent the chance F1-macro level, and the dotted lines indicate the baseline (ceiling) F1-macro obtained without frequency filtering.
7.4 Distribution shift across splits in the phoneme task
To explain the holdout generalization gap observed in the phoneme task, we quantify scale mismatches between partitions by computing the per-window root-mean-square (RMS) energy across channels and time as in Equation 3. The $x$ is the MEG window (C=306C{=}306 channels, T=125T{=}125 samples). For validation and test, windows were read with the standard competition preprocessing and standardization enabled. The holdout was read as stored (already standardized on disk). We then formed outline-only histograms using the same bin edges and density normalization across splits.
@@ 7.4 Distribution shift across splits in the phoneme task
7.5 Phonetic-feature decoding as a linguistically grounded probe
@@
As an additional analysis to probe linguistically grounded representations for MEG decoding, we trained separate binary classifiers targeting core phonetic features related to manner of articulation and voicing without sample averaging. Each model was adapted from the Phoneme Classification configuration described in Subsection 4.3, using the same backbone and training hyperparameters, but modified for binary output. Notably, we replaced the 39-class head with a single-logit output, used a binary cross-entropy loss with label smoothing and automatic positive-class weighting following Equation 1 to address class imbalance (based on the Speech model in Subsection 4.2).
This approach follows the ideas proposed by the competition authors in their blog on language-inspired phoneme classification [kwon2025], where phonetic features are suggested to handle dataset imbalance and improve infrequent phoneme decoding.
Based on [mesgarani2014], we defined six phonetic features based on manner of articulation: Plosive, Fricative, Affricate, Nasal, Liquid, Glide, and Voicing that distinguishes voiced from voiceless phonemes, including vowels and consonants. Table 4 reports preliminary decoding results using F1-macro, positive-class F1, balanced accuracy, and AUROC (mean $\pm$ std across ten seeds).
Following Subsection 7.1, statistical significance is evaluated with a one-sample Wilcoxon signed-rank test against chance ($0.5$ for F1-macro).
@@
Among these features, Voicing showed the strongest decoding signal, approaching 58% F1-macro (significant with $W=55.0$, $p=0.001$). Despite its similarity in label balance to the Speech Task, scores are not on par, implying that decoding some fine-grained speech features may not be straightforward. Manner-based features such as Plosive and Fricative were moderately decodable (both significant with $W=55.0$, $p=0.001$), while Affricate, a composite and under-represented category (0.97%), remained at chance (not significant with $W=0.0$, $p=1.0$). Attempts to address the affricate issue through ensembling of Plosive and Fricative models, transfer learning from these features, and larger effective batch sizes (via gradient accumulation) did not improve the convergence. This supports the view that decomposing phonemes into binary linguistic features alone is insufficient to overcome the data scarcity, as some linguistic features may also be rare enough not to be easily decodable, like affricates here.
@@
Overall, these experiments suggest that feature-based representations are promising for improving interpretability and possibly scaling, but further research is needed to address under-represented classes and explore multitask formulations that jointly learn shared articulatory subspaces.
8 Discussion and conclusion
The Conformer architecture proves suitable to decode the spatiotemporal nature of MEG signals, combining convolutional blocks that capture local temporal structure with self-attention for long-range dependencies. Using very similar models across both tasks, we achieved top-10 leaderboard performance, with minor input preprocessing differences between Speech and Phoneme decoding. In particular, our approach won the Phoneme Classification track in the LibriBrain Competition 2025. Instance-level normalization is indispensable to mitigate distributional shifts between training and holdout splits in the Phoneme Task, substantially improving generalization. The additional analyses in Section 7 suggest that this robustness is closely tied to distribution shift across splits. Moreover, our frequency-band analysis shows that both speech and phoneme decoding draw most of their discriminative power from low-frequency components (Delta–Beta), while Gamma and HGA contribute little in isolation. This suggests that, in the LibriBrain setting, effective MEG decoding benefits from integrating information across slow cortical dynamics, and high-frequency activity contribution is insignificant. Furthermore, a simple SpecAugment-based [park2019] augmentation was effective for speech detection, but showed limited benefit for the more complex, imbalanced phoneme classification task. On the other hand, as LibriBrain is single-subject, cross-subject generalization remains an open challenge for future work.
Looking ahead, adapting speech-model architectures for MEG decoding could open the way toward end-to-end speech reconstruction using sequential objectives such as CTC or seq2seq heads, and even enable speech synthesis. Another promising direction is the use of linguistically grounded feature-based representations [kwon2025], which may help tackle data imbalance and improve interpretability by decomposing phoneme classification into binary articulatory features (see Subsection 7.5 for the challenges found). Finally, our scaling experiments (Subsection 7.2) indicate that while speech decoding performance starts to saturate, phoneme classification continues to improve with additional data, suggesting further gains could be achieved as larger MEG datasets become available.